# Explainable Benchmarking for Iterative Optimization Heuristics

### 31 Jan 2024 v1

@@ Abstract.
Benchmarking heuristic algorithms is vital to understand under which conditions and on what kind of problems certain algorithms perform well. In most current research into heuristic optimization algorithms, only a very limited number of scenarios, algorithm configurations and hyper-parameter settings are explored, leading to incomplete and often biased insights and results. This paper presents a novel approach we call explainable benchmarking. Introducing the IOH-Xplainer software framework, for analyzing and understanding the performance of various optimization algorithms and the impact of their different components and hyper-parameters. We showcase the framework in the context of two modular optimization frameworks. Through this framework, we examine the impact of different algorithmic components and configurations, offering insights into their performance across diverse scenarios. We provide a systematic method for evaluating and interpreting the behaviour and efficiency of iterative optimization heuristics in a more transparent and comprehensible manner, allowing for better benchmarking and algorithm design.
@@
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXjournal: TELOjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Artificial intelligenceccs: Theory of computation Continuous optimizationccs: Theory of computation Bio-inspired optimizationccs: Theory of computation Design and analysis of algorithms
@@ 1. Introduction
The rapid development of new algorithmic ideas and the modification of existing algorithms in optimization heuristics (baeck2023evolutionary, ) poses a challenge in understanding their overall impact. Traditional benchmarking methods are often used to evaluate algorithms in isolation, with a single algorithm configuration (hyper-parameter setting) or with a limited set of a few variations against a limited set of state-of-the-art algorithms, leading to limited insights into their comparative performance and practical applicability. This study addresses these challenges by employing modular optimization approaches and explainable AI techniques in order to derive insights into the algorithmic behaviour of a large set of algorithm components (modules) and their hyper-parameters. Modular optimization frameworks allow for the comparison of various modifications on a core algorithm, facilitating a deeper understanding of each component’s influence on the algorithm’s performance in different scenarios. There is already a wide variety of modular algorithm frameworks available, but their application for explicit explainability of the various algorithmic components and settings has been relatively unexplored. This paper aims to bridge this gap by providing a comprehensive framework for explainable benchmarking in iterative optimization heuristics and by providing a software library (IOH-Xplainer) to facilitate researchers to use the proposed framework.
@@ 2.1. Modular Optimization Heuristics
While new algorithmic ideas or modifications of existing algorithms are proposed constantly, it is generally hard to understand the impact of these changes within a wider context. Algorithms are often benchmarked in isolation, and compared to other implementations of the same algorithm (kononova2023tiobr, ), which limits the interpretation of the results. Modular approaches to optimization algorithms resolve some of these concerns by enabling comparisons of a wide variety of modifications on the same core algorithm implementation. Modularization is achieved in such algorithms via maintaining multiple options for each algorithm operator (module) and making option choices per module in a fully independent manner since options of all modules are ensured to be compatible with each other. While modular algorithms have been around for decades, e.g. in the paradisEO framework (paradiseo22, ), and many algorithm families have been modularized (psox, ; lopez2010automatic, ), their use for the explicit purpose of explainability has been relatively limited.
@@ 2.2. Explainable AI for EC
Explainable Artificial Intelligence (XAI) has emerged as a critical field in the landscape of AI and machine learning, addressing the growing need for transparency, trustworthiness, and interpretability of AI models. As AI systems increasingly influence various aspects of daily life and critical decision-making processes, understanding how these systems arrive at their decisions has become paramount. This understanding is vital not only for the users and developers of AI systems but also for regulators and stakeholders who require assurance that AI systems are fair, safe, and aligned with ethical standards.
@@
In the context of optimization algorithms, particularly in evolutionary computation and metaheuristics, XAI plays a vital role in deciphering the mechanisms behind the algorithms’ search processes and behaviours. These algorithms often involve complex interactions and adaptations, and XAI can aid in demystifying these processes. By applying XAI techniques, researchers and practitioners can gain insights into how different components of an algorithm contribute to its overall performance, how solutions evolve over time, and what factors influence the convergence towards optimal solutions.
@@
XAI can illuminate the internal workings of algorithm components, such as selection, crossover, and mutation in genetic algorithms, or the update rules in swarm intelligence algorithms. Understanding these components through the lens of XAI can lead to the development of more efficient and effective algorithms. Furthermore, by analyzing the search behaviour of these algorithms, XAI can help identify strengths and weaknesses in different problem domains, guide the tuning of algorithm parameters, and inspire new algorithmic innovations.
@@
The field of XAI meets the field of EC in two directions, using XAI to enhance and better understand EC algorithms, and using EC to advance XAI (Zhou2024, ). For example, the approach GP-tSNE (lensen2020genetic, ) adapts the classical t-SNE algorithm, which is used for visualizing data sets, with a Genetic Programming approach to provide an interpretable mapping from the original data to the embedded points. In (evans2019s, ) the authors propose a method using multi-objective genetic programming to construct decision trees that represent any black-box classifier while being as much interpretable as possible. The proposed method aims to simultaneously maximize the ability of the tree to reconstruct the predictions of a black-box model and also maximize interpretability by minimizing the complexity of tree. The authors use a modified version of NSGA-II (deb2002fast, ). In the work conducted by Ferreira et al. (ferreira2020applying, ), a novel approach named Genetic Programming Explainer (GPX) is introduced. This method, rooted in Genetic Programming (GP), is designed to develop a localized explanatory model tailored for a specific input instance. Mirroring the strategy employed by LIME, GPX starts by selecting a target input that needs explanation. Around this chosen input, it generates a cluster of closely related data points. The key differentiation of GPX from LIME lies in its use of GP for the creation of symbolic expression trees. These trees are evolved through GP to effectively represent the functioning of the underlying complex, black-box model within the vicinity of the sampled data points. This approach allows GPX to produce more nuanced and expressive local explanations compared to the linear models typically used in LIME.
@@
  • Which hyper-parameters are important to tune?
@@ 3. Explainable Benchmarking
The field of benchmarking for black-box optimization has matured a lot over the last decades, with well respected benchmark suites such as the many versions of BBOB in the COCO platform (hansen2021coco, ), the CEC function suites (cec2017, ) and the wide variety of benchmarks within Nevergrad (nevergrad, ), and tools and analysis platforms such as IOH-Analyzer (wang2022iohanalyzer, ). Nevertheless, most of the research papers published to date which introduce new optimization heuristics contain only a very limited exploration of the capabilities and limitations of the proposed approach. Often these experimental setups are also biased towards the strengths of the newly proposed algorithm (component) to increase the chance of paper acceptance. Since we already know that one overall best heuristic does not exist (NFLT, ), it is imperative that we not only aim to develop and research optimizers that improve the state of the art (on a limited and often biased experimental setup) but also that we aim to understand the search behaviour (Mitran2023, ) and characteristics of algorithms, their inner components and their hyper-parameters. Once we have a clear understanding of how these components and their parameters influence the search (or once an AI model can capture this information), we can develop tailor-made optimisation algorithms with good performance generalisable over specific domains or landscape features.
When introducing a new algorithm or a new algorithm component, it is much more useful to know on what kind of function landscapes this new contribution has either a positive or negative influence on the performance, how much it interacts with other parameters of the algorithm configuration, and how much the component’s hyper-parameters require tuning depending on the function landscapes. This and other behaviour analyses give new insights for the community to work with. Instead, most existing works only show that the novel proposed component improves the base-algorithm (with just one or a few specific configurations) on average. Even worse, this improvement is most of the time only marginal, and it is often unclear what the underlying reasons for the improvement are, such as on what kind of function landscape, dimensionality or other landscape feature it works better or worse.
In this work, we propose a new framework for experimentation and analysis of black-box iterative optimization heuristics, we call Explainable Benchmarking, including a software framework implementation called IOH-Xplainer [^1] https://github.com/Basvanstein/ioh-explainer. In this section, we will first discuss the past and current practices of benchmarking heuristics together with their pros and cons, we will then introduce the proposed framework in detail and show how to use the framework by applying it on two large groups of modular heuristics.
@@ 3.1.1. Visual test
Over time, this ’record-average-plot-observe’ methodology persisted and evolved into reporting average performance values in time for a selection of handpicked competitor algorithms, on functions from slowly emerging benchmarking suites which are subsequently visually analyzed and discussed in the body of the paper, e.g., (kaveh2016, ). Similar approach of simple comparisons persists also in other closely related parts of the optimization community such as quality-diversity papers, e.g., (fontaine2023, ). A marginally more deep approach consists of showing a multitude of convergence plots on the entire benchmark suite and visually identifying differences in performance, as encouraged in, e.g., workshops for individual benchmarking suites like  (espinoza2023, ) from ACM GECCO Workshop on Black Box Optimization Benchmarking. Making meaningful conclusions from this multitude of plots requires significant experience. Sometimes, such plots can be enriched with statistical testing on algorithm rankings (e.g., Friedman, Nemenyi tests (mengqi2022, )). In other cases, the authors still opt for showing performance results in the table form with only mean, standard deviation, best and worst values, significantly limiting the interpretability of results.
@@ 3.1.2. Benchmarking Function Groups
The idea of identifying well-performing algorithms depending on the high-level properties of the objective functions such as multimodality, presence of global structure, separability, etc., has found its way to the setup of most benchmarking suites to test different behaviours, thus, prescribing a methodology of analyzing algorithm performance per function group, e.g.,  (pierezan2018, ; Boks2021, ). Such grouping of results can be further based on estimated problem properties (mersmann2011exploratory, ; munoz2015alg, ).
@@ 3.1.3. Domain Specific Benchmarks
Optimization algorithms are typically evaluated using now-de-facto standardised benchmark suites to discern variations in their performance. Nonetheless, the extent to which benchmark functions accurately reflect the characteristics of real-world issues remains uncertain leading to potential overfitting in algorithm choices (vanderblom2023, ). Therefore, some researchers advocate in favour of designing specialized algorithms that work well on a specific problem domain, e.g, structural optimization of vehicles (kohira2018, ), hydrothermal emission optimisation (glotic2015, ), underwater glider path planning (zamuda2019success, ). Such setup is also possible for multiobjective optimization (Tanabe2020, ).
3.2. IOH-Xplainer
@@
To analyze, compare and explainer behaviour in optimization heuristics, we propose IOH-Xplainer (Figure 1), a software package following the proposed Explainable Benchmarking framework. The framework is utilizing XAI techniques in order to automatically extract meaningful visualisations and statistics from large empirical studies, allowing the benchmarking of thousands, even millions of algorithm configurations.
@@
Figure 1. The proposed IOH-Xplainer framework, for the automatic and explainable analysis of algorithm components and hyperparameters.
@@
The proposed framework works with any (modular) algorithm, and also with collections of algorithms. By using a flexible configuration space definition (configspace, ), similar to the configuration spaces used by SMAC (lindauer2022smac3, ), we support both continuous, integer and categorical parameters. Hyper-parameter dependencies can also be defined and taken into account. IOH-Xplainer uses either a full exhaustive grid for running the experiments, or a random design with a user-defined number of algorithm configurations. Once the configuration space of algorithms and the budget to evaluate this space has been defined, each algorithm configuration (instance) is evaluated for a fixed evaluation budget and a set of dimensions on a benchmark suite. In this work, we used the well-known and popular black-box optimization benchmark (BBOB) with $24$ noiseless functions (bbobfunctions, ) originally introduced as part of the COCO framework (hansen2021coco, ). In general, the IOH-xplainer framework can be used with any benchmark. From the (parallel) benchmark runs, we collect the performance measure, which can be fixed-budget, fixed target or any-time, as is the case here. Once each instance of the experiment is finished, the explainer part of IOH-xplainer uses several explainable AI techniques to visualize the contribution of each of the algorithm’s hyper-parameters and components. First, for each function in the benchmark suite, a deep Xg-boost regression model is trained on the algorithm configuration and performance data for that function. Once the regression model is fitted to the data, a Shapley approximation technique, TreeSHAP (lundberg2020local2global, ) is utilized to extract the SHAP values for each hyper-parameter and algorithm component. These SHAP values represent the marginal contribution of the algorithm components and parameters towards the AUC of each run. When a full grid over the configuration space of algorithm instances is performed, it would even be possible to calculate the exact SHAP values from the data, however, in practice, this is infeasible due to the high computational cost. All the approximated SHAP values are aggregated per function and visualized using a swarm plot, allowing algorithm designers and practitioners to inspect the influence of each of the hyper-parameters per function.
@@
While the proposed framework uses SHAP (NIPS2017_7062, ) as local explainer, in principle any model-agnostic feature attribution XAI method can be used, such as LIME (ribeiro2016should, ) or others. In addition, global explainers can be used such as global sensitivity analysis methods (GSApaper, ; GSAcode, ; sobol1993sensitivity, ; hutter2014efficient, ) (Morris, Sobol, functional ANOVA, etc.) to understand the global first and second order sensitivity scores of each module and hyper-parameter. See our Github repository for the sensitivity analysis reports of our data.
@@
The proposed explainable benchmarking pipeline is setup in a modular fashion and allows for easy expansion in both analysis techniques as well as new algorithms and algorithms components. To use IOH-Xplainer, first a data collection step such as a large grid or large random sample of algorithm configurations has to be performed. The IOH-Xplainer toolbox runs these configurations on a benchmark suite such as BBOB for a pre-defined number of functions, instances and random seeds. Optionally the full convergence data can be logged using IOH-Analyser, otherwise only the final performance measure from each run is being logged to save storage space. Once the benchmark data has been collected, the data can be either combined with existing data collections from modCMA or modDE, or processed individually. IOH-Xplainer automatically extracts the SHAP values and creates a LaTeX report with extensive plots and tables for an easy comparison of the outcome. The toolbox can also utilize the BIAS Toolbox to perform additional checks on the single and average best configurations.

### 23 Feb 2024 v2

@@ Abstract.
Benchmarking heuristic algorithms is vital to understand under which conditions and on what kind of problems certain algorithms perform well. In most current research into heuristic optimization algorithms, only a very limited number of scenarios, algorithm configurations and hyper-parameter settings are explored, leading to incomplete and often biased insights and results. This paper presents a novel approach we call explainable benchmarking. Introducing the IOHxplainer software framework, for analyzing and understanding the performance of various optimization algorithms and the impact of their different components and hyperparameters. We showcase the framework in the context of two modular optimization frameworks. Through this framework, we examine the impact of different algorithmic components and configurations, offering insights into their performance across diverse scenarios. We provide a systematic method for evaluating and interpreting the behaviour and efficiency of iterative optimization heuristics in a more transparent and comprehensible manner, allowing for better benchmarking and algorithm design.
@@
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TELOjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Artificial intelligenceccs: Theory of computation Continuous optimizationccs: Theory of computation Bio-inspired optimizationccs: Theory of computation Design and analysis of algorithms
@@ 1. Introduction
The rapid development of new algorithmic ideas and the modification of existing algorithms in optimization heuristics (baeck2023evolutionary, ) poses a challenge in understanding their overall impact. Traditional benchmarking methods are often used to evaluate algorithms in isolation, with a single algorithm configuration (hyper-parameter setting) or with a limited set of a few variations against a limited set of state-of-the-art algorithms, leading to limited insights into their comparative performance and practical applicability. This study addresses these challenges by employing modular optimization approaches and explainable AI techniques in order to derive insights into the algorithmic behaviour of a large set of algorithm components (modules) and their hyperparameters. Modular optimization frameworks allow for the comparison of various modifications on a core algorithm, facilitating a deeper understanding of each component’s influence on the algorithm’s performance in different scenarios. There is already a wide variety of modular algorithm frameworks available, but their application for explicit explainability of the various algorithmic components and settings has been relatively unexplored. This paper aims to bridge this gap by providing a comprehensive framework for explainable benchmarking in iterative optimization heuristics and by providing a software library (IOHxplainer) to facilitate researchers to use the proposed framework.
@@ 2.1. Modular Optimization Heuristics
While new algorithmic ideas or modifications of existing algorithms are proposed constantly, it is generally hard to understand the impact of these changes within a wider context. Algorithms are often benchmarked in isolation and compared to other implementations of the same algorithm (kononova2023tiobr, ), which limits the interpretation of the results. Modular approaches to optimization algorithms resolve some of these concerns by enabling comparisons of a wide variety of modifications on the same core algorithm implementation. Modularization is achieved in such algorithms via maintaining multiple options for each algorithm operator (module) and making option choices per module in a fully independent manner since options of all modules are ensured to be compatible with each other. While modular algorithms have been around for decades, e.g. in the paradisEO framework (paradiseo22, ) and many algorithm families have been modularized (psox, ; lopez2010automatic, ), their use for the explicit purpose of explainability has been relatively limited.
@@ 2.2. Explainable AI for EC
Explainable Artificial Intelligence (XAI) has emerged as a critical field in the landscape of AI and machine learning, addressing the growing need for transparency, trustworthiness and interpretability of AI models. As AI systems increasingly influence various aspects of daily life and critical decision-making processes, understanding how these systems arrive at their decisions has become paramount. This understanding is vital not only for the users and developers of AI systems but also for regulators and stakeholders who require assurance that AI systems are fair, safe and aligned with ethical standards.
@@
In the context of optimization algorithms, particularly in evolutionary computation and metaheuristics, XAI plays a vital role in deciphering the mechanisms behind the algorithms’ search processes and behaviours. These algorithms often involve complex interactions and adaptations, and XAI can aid in demystifying these processes. By applying XAI techniques, researchers and practitioners can gain insights into how different components of an algorithm contribute to its overall performance, how solutions evolve over time and what factors influence the convergence towards optimal solutions.
@@
XAI can illuminate the internal workings of algorithm components, such as selection, crossover and mutation in genetic algorithms, or the update rules in swarm intelligence algorithms. Understanding these components through the lens of XAI can lead to the development of more efficient and effective algorithms. Furthermore, by analyzing the search behaviour of these algorithms, XAI can help identify strengths and weaknesses in different problem domains, guide the tuning of algorithm parameters and inspire new algorithmic innovations.
@@
The field of XAI meets the field of EC in two directions, using XAI to enhance and better understand EC algorithms, and using EC to advance XAI (Zhou2024, ). For example, the approach GP-tSNE (lensen2020genetic, ) adapts the classical t-SNE algorithm, which is used for visualizing data sets, with a Genetic Programming approach to provide an interpretable mapping from the original data to the embedded points. In (evans2019s, ) the authors propose a method using multi-objective genetic programming to construct decision trees that represent any black-box classifier while being as much interpretable as possible. The proposed method aims to simultaneously maximize the ability of the tree to reconstruct the predictions of a black-box model and also maximize interpretability by minimizing the complexity of trees. The authors use a modified version of NSGA-II (deb2002fast, ). In the work conducted by Ferreira et al. (ferreira2020applying, ), a novel approach named Genetic Programming Explainer (GPX) is introduced. This method, rooted in Genetic Programming (GP), is designed to develop a localized explanatory model tailored for a specific input instance. Mirroring the strategy employed by LIME, GPX starts by selecting a target input that needs explanation. Around this chosen input, it generates a cluster of closely related data points. The key differentiation of GPX from LIME lies in its use of GP for the creation of symbolic expression trees. These trees are evolved through GP to effectively represent the functioning of the underlying complex, black-box model within the vicinity of the sampled data points. This approach allows GPX to produce more nuanced and expressive local explanations compared to the linear models typically used in LIME.
@@
  • Which hyperparameters are important to tune?
@@ 3. Explainable Benchmarking
The field of benchmarking for black-box optimization has matured a lot over the last decades, with well respected benchmark suites such as the many versions of BBOB in the COCO platform (hansen2021coco, ), the CEC function suites (cec2017, ) and the wide variety of benchmarks within Nevergrad (nevergrad, ), and tools and analysis platforms such as IOHanalyzer (wang2022iohanalyzer, ). Nevertheless, most of the research papers published to date which introduce new optimization heuristics contain only a very limited exploration of the capabilities and limitations of the proposed approach. Often these experimental setups are also biased towards the strengths of the newly proposed algorithm (component) to increase the chance of paper acceptance. Since we already know that one overall best heuristic does not exist (NFLT, ), it is imperative that we not only aim to develop and research optimizers that improve the state of the art (on a limited and often biased experimental setup) but also that we aim to understand the search behaviour (Mitran2023, ) and characteristics of algorithms, their inner components and their hyperparameters. Once we have a clear understanding of how these components and their parameters influence the search (or once an AI model can capture this information), we can develop tailor-made optimisation algorithms with good performance generalisable over specific domains or landscape features.
When introducing a new algorithm or a new algorithm component, it is essential to know on what kind of function landscapes this new contribution has either a positive or negative influence on the performance, how much it interacts with other parameters of the algorithm configuration and how much the component’s hyperparameters require tuning depending on the function landscapes. This and other behaviour analyses give new insights for the community to work with. Instead, most existing works only show that the novel proposed component improves the base-algorithm (with just one or a few specific configurations) on average. Even worse, this improvement is most of the time only marginal and it is often unclear what the underlying reasons for the improvement are, such as on what kind of function landscape, dimensionality or other landscape feature it works better or worse.
In this work, we propose a new framework for experimentation and analysis of black-box iterative optimization heuristics, which we call Explainable Benchmarking, including a software framework implementation called IOHxplainer (code_xplainer_niki_van_stein_2024_10568760, ). In this section, we will first discuss the past and current practices of benchmarking heuristics together with their pros and cons, we will then introduce the proposed framework in detail and show how to use the framework by applying it on two large groups of modular heuristics.
@@ 3.1.1. Visual test
Over time, this ‘record-average-plot-observe’ methodology persisted and evolved into reporting average performance values in time for a selection of handpicked competitor algorithms, on functions from slowly emerging benchmarking suites which are subsequently visually analyzed and discussed in the body of the paper, e.g., (kaveh2016, ). A similar approach of simple comparisons persists also in other closely related parts of the optimization community such as papers on quality-diversity algorithms, e.g., (fontaine2023, ). A marginally deeper approach consists of showing a multitude of convergence plots on the entire benchmark suite and visually identifying differences in performance, as encouraged in, e.g., workshops for individual benchmarking suites like  (espinoza2023, ) from the recurring ACM GECCO Workshop on Black Box Optimization Benchmarking. Making meaningful conclusions from this multitude of plots requires attention and significant experience. Sometimes, such plots can be enriched with statistical testing on algorithm rankings (e.g., Friedman, Nemenyi tests (mengqi2022, )). In other cases, the authors still opt for showing performance results in the table form with only mean, standard deviation, best and worst values, significantly limiting the interpretability of their results.
@@ 3.1.2. Benchmarking Function Groups
The idea of identifying well-performing algorithms depending on the high-level properties of the objective functions such as multimodality, presence of global structure, separability, etc., has found its way to the setup of most benchmarking suites to test different behaviours, thus, prescribing a methodology of analyzing algorithm performance per function group, e.g.,  (pierezan2018, ; Boks2021, ). Such grouping of results can be further based on estimated problem properties like, e.g., Exploratory Landscape Analysis (ELA) features (mersmann2011exploratory, ; munoz2015alg, ).
@@ 3.1.3. Domain Specific Benchmarks
Optimization algorithms are typically evaluated using the now-de-facto standardised benchmark suites to discern variations in their performance. Nonetheless, the extent to which such benchmark functions accurately reflect the characteristics of the real-world problems remains uncertain leading to a potential overfitting in algorithmic choices (vanderblom2023, ). Therefore, some researchers advocate in favour of designing specialized algorithms that work well on a specific problem domain, e.g, structural optimization of vehicles (kohira2018, ), hydrothermal emission optimisation (glotic2015, ) and underwater glider path planning (zamuda2019success, ). Such setup is also possible for multiobjective optimization (Tanabe2020, ).
3.1.4. Modern benchmarking practices that lack immediate explainability
Despite further advancements in benchmarking methodologies (bartzbeielstein2020benchmarking, ), current practices still often overlook the necessity of ablation studies and fail to investigate the influence of parameters on algorithm performance thoroughly. This oversight can lead to an incomplete understanding of how and why certain algorithms perform better under specific conditions. A common approach in modern benchmarking involves aggregating data from multiple runs on multiple functions (but not over dimensionalities) (hansen2021coco, ). This method assumes a uniform probability distribution over a set of problems, which simplifies the analysis but may not accurately reflect real-world scenarios where problem frequencies and distributions can significantly vary. Furthermore, the reliance on absolute runtime distributions and performance profiles as universally comparable metrics across publications is problematic (hansen2021coco, ). While these metrics provide valuable insights, they often lack context regarding the scalability and adaptability of algorithms to different problem dimensions or target precisions. To address these shortcomings, it is crucial to include evaluations on the scalability of solvers across various target precisions and problem dimensions (hansen2021coco, ). Such analyses can offer more nuanced insights into the algorithm’s performance, highlighting its strengths and limitations in a more granular and explainable manner. It is furthermore recommended to limit claims of algorithm or parameter suitability strictly to the tested problem instances and to exercise caution when generalizing performance across different problem instances or classes. Such generalizations could lead to inaccuracies and overestimations of algorithmic versatility (Brownlee2007, ).
A further step in the direction of extending the explainability of benchmarking results has been taken in (vermetten2024largescale, ), where the authors outline the methodology of exploring complementarity in algorithm performances to aid automated data-driven algorithm selection and switching, assess algorithm parameterisations and implementations for completeness and tunability.
3.2. IOHxplainer
@@
To overcome some of the limitations listed above and to better analyze, compare and explain behaviour in optimization heuristics, we propose IOHxplainer (Figure 1), a software package following the proposed Explainable Benchmarking framework. The framework is utilizing XAI techniques in order to automatically extract meaningful visualisations and statistics from large empirical studies, allowing the benchmarking of thousands, even millions of algorithm configurations.
@@
Figure 1. The proposed IOHxplainer framework, for the automatic and explainable analysis of algorithm components and
hyperparameters.
@@
The proposed framework works with any (modular) algorithm and with (large) collections of algorithms. By using a flexible configuration space definition (configspace, ), similar to the configuration spaces used by SMAC (lindauer2022smac3, ), we support both continuous, integer and categorical parameters. Hyper-parameter dependencies can also be defined and taken into account. IOHxplainer uses either a full exhaustive grid for running the experiments, or a random design with a user-defined number of algorithm configurations. Once the configuration space of algorithms and the budget to evaluate this space has been defined, each algorithm configuration (instance) is evaluated for a fixed evaluation budget and a set of dimensions on a benchmark suite. In this work, we used the well-known and popular black-box optimization benchmark (BBOB) with $24$ noiseless functions (bbobfunctions, ) originally introduced as part of the COCO framework (hansen2021coco, ). In general, the IOHxplainer framework can be used with any benchmark. From the (parallel) benchmark runs, we collect the performance measure, which can be fixed-budget, fixed target or any-time, as is the case here. Once each instance of the experiment is finished, the explainer part of IOHxplainer uses several explainable AI techniques to visualize the contribution of each of the algorithm’s hyperparameters and components. First, for each function in the benchmark suite, a deep Xg-boost regression model is trained on the algorithm configuration and performance data for that function. Once the regression model is fitted to the data, a Shapley approximation technique, TreeSHAP (lundberg2020local2global, ) is utilized to extract the SHAP values for each hyper-parameter and algorithm component. These SHAP values represent the marginal contribution of the algorithm components and parameters towards the any-time performance measure of each run. When a full grid over the configuration space of algorithm instances is performed, it would even be possible to calculate the exact SHAP values from the data, however, in practice, this is infeasible due to the high computational cost. All the approximated SHAP values are aggregated per function and visualized using a swarm plot, allowing algorithm designers and practitioners to inspect the influence of each of the hyperparameters per function.
@@
While the proposed framework uses SHAP (NIPS2017_7062, ) as local explainer, in principle any model-agnostic feature attribution XAI method can be used, such as LIME (ribeiro2016should, ) or others. In addition, global explainers can be used such as global sensitivity analysis methods (GSApaper, ; GSAcode, ; sobol1993sensitivity, ; hutter2014efficient, ) (Morris, Sobol, functional ANOVA, etc.) to understand the global first and second order sensitivity scores of each module and hyper-parameter. See our Github repository (code_xplainer_niki_van_stein_2024_10568760, ) for the sensitivity analysis reports of our data.
@@
The proposed explainable benchmarking pipeline is setup in a modular fashion and allows for easy expansion in both analysis techniques as well as new algorithms and algorithm components. To use IOHxplainer, first a data collection step such as a large grid or large random sample of algorithm configurations has to be performed. The IOHxplainer toolbox runs these configurations on a benchmark suite such as BBOB for a pre-defined number of functions, instances and random seeds. Optionally the full convergence data can be logged using IOHanalyzer, otherwise only the final performance measure from each run is being logged to save storage space. Once the benchmark data has been collected, the data can be either combined with existing data collections from modCMA or modDE, or processed individually. IOHxplainer automatically extracts the SHAP values and creates a LaTeX report with extensive plots and tables for an easy comparison of the outcome. The toolbox can also utilize the BIAS Toolbox to perform additional checks on the single and average best configurations.